For each of the following relations, determine whether they are reflexive, symmetric, anti- symmetric, and/or transitive, and give a brief justification for each property. a) R ⊆ Z × Z where xRy iff x = y 2 b) The empty relation: R ⊆ A × A, where A is a non-empty set and R = ∅.
1
Expert's answer
2021-10-25T18:04:25-0400

For each of the following relations, let us determine whether they are reflexive, symmetric, anti- symmetric, and/or transitive.


a) "R ⊆ \Z × \Z" where "xRy" iff "x = y^ 2"

Since "2^2=4\ne 2," we conclude that the pair "(2,2)\notin R," and hence "R" is not a reflexive relation.


Since "4=2^2," we conclude "(4,2)\in R." On the other hand, "2\ne 4^2," and hence "(2,4)\notin R." It follows that the relation is not symmetric.


If "(x,y)\in R" and "(y,x)\in R," then "x=y^2" and "y=x^2." Therefore, "x=x^4," and hence "x=0" or "x=1." If "x=0," then "y=0." If "x=1," then "y=1." It follows that If "(x,y)\in R" and "(y,x)\in R,"

then "x=y=0" or "x=y=1," and hence the relation is antisymmetric.


Taking into account that "(16,4)\in R" and "(4,2)\in R," but "(16,2)\notin R," we conclude that the relation "R" is not transitive.


b) The empty relation: "R \u2286 A \u00d7 A" , where "A" is a non-empty set and "R = \u2205".

Let "a\in A." Since "(a,a)\notin\emptyset= R," we conclude that the relation "R" is not reflexive.


Since the statement "[(x,y)\in\emptyset= R]" is false, the implication "[(x,y)\in R]\to[(y,x)\in R]" is true, and hence this relation is symmetric.


Taking into account that the statement "[(x,y)\in R\text{ and }(y,x)\in R]" is false, the implication "[(x,y)\in R\text{ and }(y,x)\in R]\to [x=y]" is true, and hence this relation is antisymmetric.


Since the statement "[(x,y)\in R\text{ and }(y,z)\in R]" is false, the implication "[(x,y)\in R\text{ and }(y,z)\in R]\to [(x,z)\in R]" is true, and hence this relation is transitive.